OpenStudy (anonymous):

how would you transform f(x) = 3 sin(4x - π) + 4 into a cosine function in the form f(x) = a cos(bx - c) + d?

OpenStudy (amistre64):

what is the relationship between sine and cosine?

OpenStudy (amistre64):

in a right triangle, with angles a and b a+b+90 = 180, b = 90-a sin(a) = cos(b) sin(a) = cos(90-a)

OpenStudy (amistre64):

recall 90 degrees is pi/2 radians which might be more prudent in this case